2x-5y+4=0 and passes through the point (1,0)

Found 2 solutions by Cromlix, ikleyn:
Answer by Cromlix(4381)

Hi there,
2x - 5y + 4 = 0
y = 2/5x + 4/5
A line parallel to the above line would have a slope of 2/5
and passing through point (1, 0) would have an equation of
y = 2/5x - 2/5 or 5y = 2x - 2
A line perpendicular to the above line would have a slope of - 5/2
and passing through point (1, 0) would have an equation of
y = - 5/2x + 5/2 or 2y = -5x + 5
Hope this helps :-)

Answer by ikleyn(52781)

.

How the problem is presented in the post, it is   D E F E C T I V E.

Please avoid posting NONSENSE to the forum.